OpenStudy (anonymous):

plzz help me to find the domain and range of f(x) =-sqrt(-5-6x-x^2)

OpenStudy (tkhunny):

Well, start with -5-6x-x^2 >= 0 or 5 + 6x + x^2 <= 0

OpenStudy (anonymous):

I got the domain D = [-5,-1] , not getting the range got y^2=4 so it should be -2<=y<=2 but answer for range is [-2,0]

OpenStudy (anonymous):

y is always negative, so it should be from [-2,0]

OpenStudy (tkhunny):

The Domain is Negative Only. You cannot get a positive number out of that. Excellent job ont he Domain.

OpenStudy (anonymous):

how do i write it ? what should be my statements after y^2=4 or -2<=y<=2

OpenStudy (anonymous):

well, there is a negative sign infront , so that makes the range always negative

OpenStudy (tkhunny):

No, that still have y > 0. Eliminate that part.

OpenStudy (anonymous):

I want to frame the answer with right wordings could you tell me what to write to arrive at Range [-2,0]

OpenStudy (anonymous):

of course, i didn't get [-2,0] right off the bat. Notice that -sqrt(-5-6x-x^2 = -sqrt[4 - (x+3)^2)] This is just a lower semi-circle whose center is at (-3,0) and with radius = 2. That is how I got [-2,0]. I just assumed you knew this

OpenStudy (anonymous):



OpenStudy (anonymous):

the lower part :DD

OpenStudy (tkhunny):

We have a lovely parabola, y = -5-6x-x^2 = -(x+3)^2 + 4 Notice, with the vertex at (-3,4), we have the range \([-\infty,4)\). Now, we're going to need to find the square root of that, so we discard any negative portion. Thus, the range is now \([0,4]\) Once we find the square root (positive, unless otherwise stated), we find that the Range is now \([0,2]\) Finally, some sadistic textbook author put a negative sign in front of this one, so we're stuck with a Range of \([-2,0]\) You just have to walk through it.

OpenStudy (tkhunny):

Whoops, that first one should have been \((-\infty,4]\).